OpenStudy (anonymous):

simplify √72-3√12+√192

OpenStudy (kuoministers):

goodmorning

OpenStudy (kuoministers):

first thing we need to do is to break down the surds so they are the same rooted number

OpenStudy (kuoministers):

e.g \[\sqrt{18}\] = \[3\sqrt{2}\]

OpenStudy (anonymous):

ok

OpenStudy (kuoministers):

so what can sqrt 72 be broken down to?

OpenStudy (anonymous):

9?

OpenStudy (kuoministers):

hint: 36 could be one factor

OpenStudy (kuoministers):

9 is good but we need to find the largest one which is 36!!

OpenStudy (anonymous):

ok!

OpenStudy (kuoministers):

\[\sqrt{72} = 6\sqrt{2}\]

OpenStudy (kuoministers):

now can you do sqrt 12?

OpenStudy (anonymous):

2 sqrt 3?

OpenStudy (kuoministers):

good!! now for sqrt 192?

OpenStudy (anonymous):

i dont kno

OpenStudy (kuoministers):

just keep trying all the \[2^{2} 3^{2}\]

OpenStudy (kuoministers):

its pretty much trial and error but hint is 64

OpenStudy (anonymous):

ok thanks

OpenStudy (kuoministers):

now can you solve the question?

OpenStudy (kuoministers):

\[6\sqrt{2} - 6\sqrt{3} + 8\sqrt{3}\]

OpenStudy (kuoministers):

Do you understand it now?

OpenStudy (anonymous):

yea thanks

OpenStudy (kuoministers):

if you think my answer helped you click best response Thanks